Give the resultant vector of A = 26.5 m 56° NW; B = 44.0 m 28° NE; and C = 31.0 m S
1
Expert's answer
2020-10-21T09:53:05-0400

Vector A:


"A=-26.5\text{ sin}56°i+26.5\text{cos}56°j=-14.8i+22j."


Vector B:


"B=44\text{ sin}28°i+44\text{ cos}28°j=38.8i+20.7j."


Vector C:


"C=-31j."

The resultant is

"R=A+B+C=(24i+11.7j)\text{ km}"